OpenStudy (anonymous):

Another independent probability question?? Algebra 2

OpenStudy (anonymous):

A two-way frequency table shows grades for students in college and students in high school.

OpenStudy (anonymous):

Based on this data, are "being in high school" and "GPA above 3.0" independent events? Yes, P(high school | GPA above 3.0) = P(high school) Yes, P(high school | GPA above 3.0) = P(GPA above 3.0) No, P(high school | GPA above 3.0) ≠ P(high school) No, P(high school | GPA above 3.0) ≠ P(GPA above 3.0)

OpenStudy (anonymous):

@satellite73 @jennaknight03

OpenStudy (anonymous):

I am really bad with this kind of problem so i can not help sorry :(

OpenStudy (anonymous):

now this one at least makes sense

OpenStudy (anonymous):

It's okay, Jenna :)

OpenStudy (anonymous):

thanks @TrixieDiVine

OpenStudy (anonymous):

the probability that you went to high school, given that your gpa is above 3.0, is \(\frac{14}{40}\) since \(40\) total had a gpa of above 3.0 and of those 14 had high school

OpenStudy (anonymous):

the total of the people who went to high school is 60, the total sample is 100 so the probability of high school is \(\frac{60}{100}=.6\)

OpenStudy (anonymous):

compare \(.6\) with \(\frac{14}{40}\) and you will see that they are not equal, so the answer is DEPENDENT

OpenStudy (anonymous):

go with No, P(high school | GPA above 3.0) ≠ P(high school)

OpenStudy (anonymous):

as always, it is C the probability that it is C is \(.9\) for any multiple choice question

OpenStudy (anonymous):

I thought that was it! :D

OpenStudy (anonymous):

is this your last one if not mention me and i will try to help